OpenStudy (anonymous):

Doing a review and need help please /: Which of the following is a solution of x2 + 2x + 8?

OpenStudy (anonymous):

\[4 + i \sqrt{7}\] \[-1 + i \sqrt{7}\] \[1+i \sqrt{7}\] \[-4 + i \sqrt{7}\]